Grade 9Algebra

a-1/a = 3
find a^2 + 1/a^2
give me solution for the above

devendra jaiswal
9 Years agoGrade 9

1 Answer

Anshika Pattnaik
9 Years ago
a-1/a=3 (given)(a-1/a)^square = 3^square (squaring both sides.)a^square +(1/a)^square - 2(a*1/a) = 9[(a-b)^square = a^square + b^square - 2a*b]a^square + (1/a) ^square -2 = 9a^square + (1/a) ^square = 9+2 = 11